OpenStudy (rileystevens):

For equation y=(1/3)^x does it have an inverse? When I try to find the inverse, should I switch 1/3 into a decimal and then do x=(0.33333)^y and solve?

OpenStudy (3mar):

May I help?

OpenStudy (3mar):

To get the inverse of any function f(x): - Solve for x, i.e separate x in one side and the other terms in the other side, included f(x). - Swap x and f(x). - The result is the inverse of the original function.

OpenStudy (rileystevens):

Yes please, I'm trying to get an assignment done and I have to find the inverse of \[y=5^{x} \] which I thought would be \[x=5^{y} \] but when I substitute x for a negative number to graph I don't think you can solve \[-5=5^{y}\]. And then for \[y=\frac{ 1 }{ 3 }^{x} \] I'm not sure if it would be \[x=\frac{ 1 }{ 3 }\] and if that would make it \[x=(0.333333..)^{y}\]

OpenStudy (3mar):

Great job! With my pleasure!

OpenStudy (3mar):

Can you follow my steps and tell me what you got?

OpenStudy (3mar):

@rileystevens

OpenStudy (rileystevens):

Perfect

OpenStudy (triciaal):

from above. " but when I substitute x for a negative number to graph I don't think you can solve −5=5y" need to solve for y = the graph is y as a function of x

OpenStudy (rileystevens):

so would it be \[\frac{ -5 }{ 5 }=\frac{ 5y }{ 5 }\] -1=y

OpenStudy (danjs):

\[\large y=5^x\] For that function, think about what happens as you start putting in larger and larger negative numbers for x. The y value becomes smaller and smaller approaching zero but never getting there. The graph is...